OpenStudy (anonymous):

Find the smaller solution for x^2−20=−8

OpenStudy (anonymous):

\[x ^{2}-20=-8 \] first subtract add 20 to both sides which is: x^2=12 then take the square root of 12 which is 2 square root 3

OpenStudy (anonymous):

*i meant add 20 to both sides

OpenStudy (anonymous):

You need to isolate x and we do this by grouping like terms together. The best way to start this is to get rid of that -20 on the left. We do this by adding 20, and what we do on the left we must do on the right. Now that we have x isolated, we need to get rid of that exponent so we square root both sides. x^2 -20 = -8 x^2 - 20 + 20 = -8 + 20 x^2 = 12 sqrt(x^2) = sqrt(12) x = sqrt(12)